OpenStudy (anonymous):

John wants to rent movies over his school vacation. The cost for the movie rental is $4.50 per movie. John starts out with $20. Write an equation for the amount of money, y, John will have after x amount of movie rentals. Answer y = 20 - 4.50x y = 4.50x - 20 y = 20 + 4.50x y = 4.50x

OpenStudy (anonymous):

@MiaKitty @UnkleRhaukus Help

OpenStudy (anonymous):

This is just like the question I helped you with earlier, but you have to combine them together, what do you think?

OpenStudy (anonymous):

Either B or C

OpenStudy (unklerhaukus):

\[y=mx+b\] where \(m\) is the rate of change and \(b\) is the initial value

OpenStudy (anonymous):

Well, it isn't B, because you don't subtract values in this equation, so?

OpenStudy (anonymous):

I have a question for you, ramiro: Let's say you buy two movies. How much money will you be left with?

OpenStudy (anonymous):

Ummm probably 0

OpenStudy (anonymous):

or at least 3 bucks or something

OpenStudy (anonymous):

Hmm... tell me this: how much would two movies cost?

OpenStudy (anonymous):

I have to go now! Anyone who can help Ramirocool can. :)

OpenStudy (anonymous):

@UnkleRhaukus I got A

OpenStudy (unklerhaukus):

note \(mx+b=b+mx\)

OpenStudy (anonymous):

@UnkleRhaukus D.

OpenStudy (anonymous):

@UnkleRhaukus : I think at Ramiro's age, formulas like that are more confusing that actually using real life examples. Granted, I'm under the impression that Ramiro doesn't care either way, so it probably doesn't matter.

OpenStudy (unklerhaukus):

if one movie costs $4.50 two movies costs 2×$4.50=

OpenStudy (anonymous):

8.50

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

Close. But do it on paper to be safe.

OpenStudy (anonymous):

@uber1337h4xx0r /yup B

OpenStudy (unklerhaukus):

\[2\times4.50=2\times4+2\times0.50=8+1=\]

OpenStudy (anonymous):

@UnkleRhaukus Sorry for the long wait C.

OpenStudy (anonymous):

Right?